The overhead reach distances of adult females are normally distributed with a mean of 200 cm and a standard deviation of 7.8 cm.

a. Find the probability that an individual distance is greater than 209.30 cm.

b. Find the probability that the mean for 25 randomly selected distances is greater than 197.80 cm.

c. Why can the normal distribution be used in part​ (b), even though the sample size does not exceed​ 30?
Answer #1

The overhead reach distances ( X ) of adult females are normally distributed

Mean = \mu = 200 cm , standard deviation = \sigma = 7.8 cm

a )

We have to find P( x > 209.30 )

P( x > 209.30 ) = 1 - P( x <= 209.30 )

Using Excel function,   =NORMDIST( x , mean , SD , 1 )

P( x <= 209.30 ) = NORMDIST( 209.30 , 200, 7.8 , 1 ) = 0.88343

So, P( x > 209.30 ) = 1 - 0.88343 = 0.11657

The probability that an individual distance is greater than 209.30 cm is 0.11657

b)

We have to find P( \bar{x} > 197.80 ) , sample size = n = 25

Mean of \bar{x} = 200

Standard deviation of \bar{x} = a/Vn = 7.8/ 25 = 1.56

P( \bar{x} > 197.80 ) = 1 - P( \bar{x} <= 197.80 )

Using Excel function,   =NORMDIST( x , mean , SD , 1 )

P( \bar{x} <= 197.80 ) = NORMDIST(197.80 , 200 , 1.56 , 1 ) = 0.079232

So, P( \bar{x} > 197.80 ) = 1 - 0.079232 = 0.92077

The probability that the mean for 25 randomly selected distances is greater than 197.80 cm is 0.92077

c )

The normal distribution can be used in part (b) because the original population has a normal distribution.